# Attention-aware Semantic Communications for Collaborative Inference

### v1

@@ VI-C Attention-aware Patch Selection
In this subsection, we delve into identifying the most effective metrics for determining patch importance and the optimal rules for patch selection. First, we examine and compare the efficacy of mean attention scores versus attention rollout. Next, we evaluate different patch selection methodologies, including top-$k$ selection, attention threshold selection, and attention-sum threshold selection. We maintain a consistent environment with DeiT-Tiny on the edge device and DeiT-Base on the server. For this analysis, we specifically focus on attention-aware patch selection, excluding considerations of entropy-aware image transmission.
Figure 11: Comparison of mean attention score and attention rollout, with the edge device employing DeiT-Tiny and the server employing DeiT-Base. The patch selection rule is the attention-sum threshold selection. (a) shows the overall trade-off between communication cost and classification accuracy, while (b) zooms in on the region of interest, focusing on areas near the server model’s classification accuracy.
Fig. 11LABEL:sub@fig:attention-measure-a reveals that both the mean attention score and attention rollout have distinct advantages. In situations demanding substantial reductions in communication cost, attention rollout is better than the mean attention score. However, as shown in Fig. 11LABEL:sub@fig:attention-measure-b, the mean attention score is a better metric for attaining accuracy comparable to that of the server. Consequently, we adopt the mean attention score for our primary experiments.
@@
Fig. 12 shows that both attention threshold selection and attention-sum threshold selection outperform the performance of the top-$k$ selection. These methods offer the capability to adjust the number of selected patches for transmission, facilitating the maintaining classification accuracy while minimizing communication costs.
@@ VI-D Entropy-aware Image Transmission
VI-E Model Complexity of Semantic Encoder
As discussed in Section IV-A, DeiT-Tiny can act more effectively as a semantic encoder than DeiT-Base in spite of its inferior classification accuracy, which is supported by Fig. 5 and Fig. 6. In our experiments, we concentrate solely on attention-aware patch selection to evaluate and compare DeiT-Tiny and DeiT-Base. Both models select only important patches based on the mean attention scores, disregarding initial inference outcomes. The classification accuracy is obtained by DeiT-Base, which processes only these selected image patches. Fig. 14 shows that DeiT-Tiny more effectively identifies the essential image patches for classification better than DeiT-Base.
Figure 14: Comparison of DeiT-Tiny and DeiT-Base as semantic encoders. The classification accuracy is obtained by DeiT-Base.
@@ VII Conclusion
We presented a novel attention-aware collaborative inference system using pre-trained ViT models. The edge device employs a compact ViT model as a semantic encoder, selecting pivotal patches that focus on objects crucial for classification. Our results confirm that this framework not only significantly lowers communication costs but also preserves accuracy comparable to that of server models. Furthermore, it offers the added benefit of diminishing the server model’s computational complexity. Extending this approach to encompass a broader range of tasks beyond classification provides an intriguing direction for future research.

### v2

@@ VI-C Attention-aware Patch Selection
In this subsection, we delve into identifying the effective attention metrics for determining patch importance and the optimal rules for patch selection. First, we examine and compare the efficacy of mean attention scores versus attention rollout. Next, we evaluate different patch selection methodologies, including top-$k$ selection, attention threshold selection, and attention-sum threshold selection. We maintain a consistent environment with DeiT-Tiny on the edge device and DeiT-Base on the server. For this analysis, we specifically focus on attention-aware patch selection, excluding considerations of entropy-aware image transmission.
Figure 11: Comparison of mean attention score and attention rollout, with the edge device employing DeiT-Tiny and the server employing DeiT-Base on the ImageNet dataset. The patch selection rule is the attention-sum threshold selection. (a) shows the overall trade-off between communication cost and classification accuracy, while (b) zooms in on the region of interest, focusing on areas near the server model’s classification accuracy. (c) shows the normalized histograms of mean attention score and attention rollout, obtained by DeiT-Tiny.
Fig. 11LABEL:sub@fig:attention-measure-a reveals that both the mean attention score and attention rollout have distinct advantages. In situations demanding substantial reductions in communication cost, attention rollout is better than the mean attention score. However, as shown in Fig. 11LABEL:sub@fig:attention-measure-b, the mean attention score is a better metric for attaining accuracy comparable to that of the server.
The theoretical advantage of the mean attention score over attention rollout is shown in Fig. 11LABEL:sub@fig:attention-measure-c, which compares the normalized histograms of attention scores for both metrics. Unlike attention rollout, which considers attention scores from multiple layers resulting in a more uniform distribution of scores, the mean attention scores concentrate on fewer patches with significantly higher relevance to the objects to be classified. This concentration of attention scores is preferred because it implies that fewer but more relevant patches can be transmitted without compromising the classification accuracy, thus aligning with our objective to minimize communication overhead while maintaining classification accuracy.
Hence, we adopt the mean attention score for our primary experiments.
@@
Fig. 12 shows that both the attention threshold selection and attention-sum threshold selection outperform the top-$k$ selection. These methods offer the capability to adjust the number of selected patches for transmission, facilitating the maintaining classification accuracy while minimizing communication costs.
The variability in informational content across patches, influenced by factors such as object size, type, and the presence of background, directly impacts their importance for accurate classification. Theoretically, the attention scores reflect each patch’s contribution to the model prediction. By setting thresholds based on cumulative attention scores (attention-sum threshold) or individual patch attention scores (attention threshold), we can dynamically control the number of transmitted patches based on their estimated relevance.
@@ VI-D Entropy-aware Image Transmission
VI-E Comparison of Models as Semantic Encoder
As discussed in Section IV-A, DeiT-Tiny can act more effectively as a semantic encoder than DeiT-Base in spite of its inferior classification accuracy. In our experiments, we concentrate solely on attention-aware patch selection to evaluate and compare DeiT-Tiny and DeiT-Base. Both models select only important patches based on the mean attention scores, disregarding initial inference outcomes. The classification accuracy is obtained by DeiT-Base, which processes only these selected image patches. Fig. 14 shows that DeiT-Tiny more effectively identifies the essential image patches for classification better than DeiT-Base.
In particular, Fig. 5 and Fig. 6 in Section IV-A show that the attention scores obtained by DeiT-Tiny concentrate more on relevant objects than those from DeiT-Base. Recent work  [32] supports this observation, showing that large ViT models often allocate high attention scores to less informative background areas. It is because the large ViT models adeptly identify patches containing minimal information, such as background areas, repurposing the corresponding patches to assimilate global image information while neglecting spatial information. While these high attention score patches may enhance the classification accuracy, they diminish the effectiveness of base models as semantic encoders.
Figure 14: Comparison of DeiT-Tiny and DeiT-Base as semantic encoders to select important patches. The classification accuracy is obtained by DeiT-Base.
VI-F Inference Latency Analysis
In this subsection, we evaluate the end-to-end inference latency of both our collaborative inference and the conventional server-based inference.
The end-to-end inference latency consists of client inference latency, server inference latency, and communication latency. Client inference latency, set as 10.99 milliseconds (ms), reflects the inference latency of DeiT-Tiny on the iPhone 12 neural engine [18]. Server inference latency is set at 8.32 ms, corresponding to the inference latency of DeiT-Base on an NVIDIA RTX 3090 GPU [22]. This latency can be eliminated if the image is not transmitted to the server via the strategy of entropy-aware image transmission. Additionally, server inference latency can be reduced in proportion to the decreased computational complexity (FLOPs) if only selected patches are transmitted via the strategy of attention-aware patch selection. The FLOPs of DeiT-Base are calculated by (9). Communication latency is calculated using the transmitted data size divided by the upload data rate. We estimate the data size of a typical image to be 147 KB, based on an image cropped to $224\times 224\times 3$ bytes. The upload data rates of 1 Mbps, 8 Mbps, and 20 Mbps are considered as in [44].
Figure 15: Comparison of inference latency for different upload data rates, 1 Mbps, 8 Mbps, and 20 Mbps.
Figure 16: Breakdown of end-to-end inference latency of the proposed collaborative inference.
[…]
@@ VII Conclusion
We presented a novel attention-aware collaborative inference framework using pre-trained ViT models. The edge device employs a lightweight ViT model as a semantic encoder, selecting pivotal patches that focus on objects crucial for classification. Our results confirm that this framework not only significantly lowers communication costs but also preserves accuracy comparable to that of server models. Furthermore, it offers the added benefit of diminishing the server model’s computational complexity. Extending this approach to encompass a broader range of tasks beyond classification provides an intriguing direction for future research.